# Know What You Don’t Know: Uncertainty Calibration of Process Reward Models

### v1

@@ E.3 Inference-time Scaling with IAS
We first present results and analysis of the proposed IAS strategy across different LLMs and PRMs. Specifically, we report performance under the best-of-$N$ framework in Tables 2, 4, and 5. To ensure statistical significance, the pass@1 rate is estimated by averaging accuracy over $64$ independent forward passes. Similarly, for the IAS approach, we report the score based on $100$ different trials for each question, based on the determined $N_{\mathrm{IAS}}$ value computed individually per question.
As observed, the IAS approach adaptively allocates the sampling budget while largely preserving overall accuracy. Ironically, the ReasonEval PRM performs poorly under the best-of-$N$ (BoN) strategy, often yielding results worse than the pass@1 rate. However, this is not a limitation of the IAS strategy itself, but rather a reflection of the ReasonEval PRM’s weakness in selecting the correct answer from among candidate responses. Notably, the budgets determined by IAS for ReasonEval are comparable to those derived from the other PRMs (See Figures 11, 12, and 13 for how effectively IAS adaptively selects $N$.). In contrast, PRMs like Qwen exhibit strong ranking capabilities despite poor initial calibration. Our calibration method proves particularly effective in such cases, as it preserves the PRM’s original strengths while enabling effective IAS and improving its reliability.
@@
Tables 3, 6, and 7 present the pass@1 accuracy for standard beam search (BS), BS with the proposed IASoK, and BS with IASoM, evaluated using the Qwen, Shepherd, and ReasonEval PRMs. For the Shepherd and ReasonEval PRMs, we conduct experiments using the MATH100 (i.e., the first $100$ examples among MATH100) and AIME2024 datasets. To ensure a fair comparison, all experiments were conducted using the calibrated PRM. The results are consistent with the observations made for the Qwen PRM, as discussed in the main text.
E.4 Ablation Studies on IAS Parameters
@@
The target probability serves as a parameter that acts as a constant multiplier on the expression $1/\log(1-p)$. We present results for values of $C$ such that $\log(1-C)$ takes on values $-0.125$, $-0.25$, $-0.5$, $-1.0$, $-2.0$, $-3.0$, $-4.0$, and $-5.0$. Figure 14 shows the budget-accuracy trade-off plots. We use $C=0.99$, as it appears to provide a reasonable balance between accuracy and computational cost.
@@
Again, as also shown in Figures 11, 12, and 13, accuracy tends to increase monotonically with higher inference budget, thus a “sweet spot” doesn’t exist. Exploring a principled metric to assess the cost-performance ratio within the inference-time scaling framework remains an interesting direction for future work.
@@
Table 4: Evaluation of the best-of-$N$ method with a fixed-$N$ strategy (BoN), compared against the proposed instance-adaptive sampling strategy (BoN+IAS), with calibrated Shepherd PRM. We report both accuracy and normalized computational cost (budget). Relative improvements over Pass@1 accuracy are highlighted in light blue.
@@
Table 5: Evaluation of the best-of-$N$ method with a fixed-$N$ strategy (BoN), compared against the proposed instance-adaptive sampling strategy (BoN+IAS), with calibrated ReasonEval PRM. We report both accuracy and normalized computational cost (budget). Relative improvements/degradations over Pass@1 accuracy are highlighted in light blue/red.
@@
Table 6: Evaluation of the beam search method with a fixed-$N$/$M$, compared against the proposed adaptive sampling strategy (IASoK and IASoM) using calibrated Shepherd PRMs. We report both accuracy and computational cost (budget), measured by the average number of LLM generations per question normalized by that of a fixed-budget BS baseline. Relative improvements over the Pass@1 accuracy are highlighted in light blue.
@@
Table 7: Evaluation of the beam search method with a fixed-$N$/$M$, compared against the proposed adaptive sampling strategy (IASoK and IASoM) using calibrated ReasonEval PRMs. We report both accuracy and computational cost (budget), measured by the average number of LLM generations per question normalized by that of a fixed-budget BS baseline. Relative improvements/degradations over Pass@1 accuracy are highlighted in light blue/red.
@@
Figure 11: IAS scales with problem difficulty in MATH500, with Qwen PRM. Average accuracy over the test points with varying difficulty levels (1: easy, 5: hard) is illustrated. Regular fixed-$N$ and our IAS approaches are indicated by dashed lines and stars, respectively, and IAS allocates more samples to harder questions.
Figure 12: IAS scales with problem difficulty in MATH500, with Shepherd PRM. Average accuracy over the test points with varying difficulty levels (1: easy, 5: hard) is illustrated. Regular fixed-$N$ and our IAS approaches are indicated by dashed lines and stars, respectively, and IAS allocates more samples to harder questions.
Figure 13: IAS scales with problem difficulty in MATH500, with ReasonEval PRM. Average accuracy over the test points with varying difficulty levels (1: easy, 5: hard) is illustrated. Regular fixed-$N$ and our IAS approaches are indicated by dashed lines and stars, respectively, and IAS allocates more samples to harder questions.

### v2

@@ E.4 Inference-time Scaling with IAS
We first present results and analysis of the proposed IAS strategy across different LLMs and PRMs. Specifically, we report performance under the best-of-$N$ framework in Tables 2, 6, and 7. To ensure statistical significance, the pass@1 rate is estimated by averaging accuracy over $64$ independent forward passes. Similarly, for the IAS approach, we report the score based on $100$ different trials for each question, based on the determined $N_{\mathrm{IAS}}$ value computed individually per question.
As observed, the IAS approach adaptively allocates the sampling budget while largely preserving overall accuracy. Ironically, the ReasonEval PRM performs poorly under the best-of-$N$ (BoN) strategy, often yielding results worse than the pass@1 rate. However, this is not a limitation of the IAS strategy itself, but rather a reflection of the ReasonEval PRM’s weakness in selecting the correct answer from among candidate responses. Notably, the budgets determined by IAS for ReasonEval are comparable to those derived from the other PRMs (See Figures 12, 13, and 14 for how effectively IAS adaptively selects $N$.). In contrast, PRMs like Qwen exhibit strong ranking capabilities despite poor initial calibration. Our calibration method proves particularly effective in such cases, as it preserves the PRM’s original strengths while enabling effective IAS and improving its reliability.
@@
Tables 3, 8, and 9 present the pass@1 accuracy for standard beam search (BS), BS with the proposed IASoK, and BS with IASoM, evaluated using the Qwen, Shepherd, and ReasonEval PRMs. For the Shepherd and ReasonEval PRMs, we conduct experiments using the MATH100 (the first $100$ examples among MATH500) and AIME2024 datasets. To ensure a fair comparison, all experiments were conducted using the calibrated PRM. The results are consistent with the observations made for the Qwen PRM, as discussed in the main text.
E.5 Ablation Studies on IAS Parameters
@@
The target probability serves as a parameter that acts as a constant multiplier on the expression $1/\log(1-p)$. We present results for values of $C$ such that $\log(1-C)$ takes on values 0.125-0.125, 0.25-0.25, 0.5-0.5, 1.0-1.0, 2.0-2.0, 3.0-3.0, 4.0-4.0, and 5.0-5.0. Figure 15 shows the budget-accuracy trade-off plots. We use $C=0.99$, as it appears to provide a reasonable balance between accuracy and computational cost.
@@
Again, as also shown in Figures 12, 13, and 14, accuracy tends to increase monotonically with higher inference budget, thus a “sweet spot” doesn’t exist. Exploring a principled metric to assess the cost-performance ratio within the inference-time scaling framework remains an interesting direction for future work.
@@
Table 6: Evaluation of the best-of-$N$ method with a fixed-NN strategy (BoN), compared against the proposed instance-adaptive sampling strategy (BoN+IAS), with calibrated Shepherd PRM. We report both accuracy and normalized computational cost (budget). Relative improvements over Pass@1 accuracy are highlighted in light blue.
@@
Table 7: Evaluation of the best-of-$N$ method with a fixed-NN strategy (BoN), compared against the proposed instance-adaptive sampling strategy (BoN+IAS), with calibrated ReasonEval PRM. We report both accuracy and normalized computational cost (budget). Relative improvements/degradations over Pass@1 accuracy are highlighted in light blue/red.
@@
Table 8: Evaluation of the beam search method with a fixed-NN/$M$, compared against the proposed adaptive sampling strategy (IASoK and IASoM) using calibrated Shepherd PRMs. We report both accuracy and computational cost (budget), measured by the average number of LLM generations per question normalized by that of a fixed-budget BS baseline. Relative improvements over the Pass@1 accuracy are highlighted in light blue.
@@
Table 9: Evaluation of the beam search method with a fixed-NN/$M$, compared against the proposed adaptive sampling strategy (IASoK and IASoM) using calibrated ReasonEval PRMs. We report both accuracy and computational cost (budget), measured by the average number of LLM generations per question normalized by that of a fixed-budget BS baseline. Relative improvements/degradations over Pass@1 accuracy are highlighted in light blue/red.
@@
Figure 12: IAS scales with problem difficulty in MATH500, with Qwen PRM. Average accuracy over the test points with varying difficulty levels (1: easy, 5: hard) is illustrated. Regular fixed-NN and our IAS approaches are indicated by dashed lines and stars, respectively, and IAS allocates more samples to harder questions.
Figure 13: IAS scales with problem difficulty in MATH500, with Shepherd PRM. Average accuracy over the test points with varying difficulty levels (1: easy, 5: hard) is illustrated. Regular fixed-NN and our IAS approaches are indicated by dashed lines and stars, respectively, and IAS allocates more samples to harder questions.
Figure 14: IAS scales with problem difficulty in MATH500, with ReasonEval PRM. Average accuracy over the test points with varying difficulty levels (1: easy, 5: hard) is illustrated. Regular fixed-NN and our IAS approaches are indicated by dashed lines and stars, respectively, and IAS allocates more samples to harder questions.